# Laura Barros        Aiko Schmeißer

### v3

@@ 4.2 Main results
We begin by discussing the effects of job loss on labor market trajectories. Figure 1 displays yearly event-study estimates based on model (3). In the year of the layoff, workers experience a sharp decline in labor earnings of around 10,000 BRL (2,028 USD), which correspondents to a 45% reduction relative to baseline earnings in the year prior to the layoff. Moreover, the probability to be formally employed at the end of the year falls by about 60 percentage points. Starting two years after the layoff, both earnings and employment start to continuously recover; yet, the recovery is only partial and does not return to pre-layoff levels. Seven years after the layoff – the longest horizon we can observe in the RAIS data – dismissed workers still earn about 5,000 BRL (1,014 USD) less per year and are 15 percentage points less likely to be formally employed compared to the control group. Appendix Figure B.1 further shows that job loss reduces subsequent job stability, as reflected in a higher frequency of job separations. These findings align with the job loss literature that has widely documented the long-term scarring effects of job displacements (e.g., Jacobson et al., 1993; Eliason and Storrie, 2006; Davis and von Wachter, 2011; Lachowska et al., 2020; Schmieder et al., 2023). [^26] Britto et al. (2022b) also show, using survey data on Brazilian workers, that accounting for income from the informal sector reduces the estimated earnings losses following layoffs by only about 10%.
@@
Effect timing. Next, we take a closer look at the dynamics in the estimated treatment effects. Given that deciding to become a candidate and preparing for an election campaign requires time and planning, we start by examining whether our results differ by proximity to the next local election. Appendix Figure B.3 shows results when splitting the sample by the layoff year (2009, 2010, or 2011). We find that the positive effects of job loss on party memberships and candidacies are not driven by one specific treatment year, suggesting that temporal proximity to the local election does not matter. This dynamic also supports the external validity of our results. For instance, there may be the concern that political responses following layoffs in 2009 could differ from other layoffs, given the proximity to the 2008 financial crisis. However, our results show similar effects for layoffs occurring in 2010 and 2011, once the crisis had already faded out in Brazil (Barbosa, 2010).
@@
A. Running for mayor B. Campaign donation Refer to caption Refer to caption C. Campaign workRefer to caption Note: The figure shows results for the effect of job loss on the likelihood of running for municipal mayor (Panel A), donating to a political campaign (Panel B), and working for a political campaign (Panel C). Information on mayoral candidates and campaign donations are obtained from the Tribunal Superior Eleitoral. The data on campaign donations is only available starting from the 2004 election cycle. Following Bazzi and Labanca (2023), campaign workers are identified in RAIS as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Panels A and B report cycle-specific event-study coefficients $\delta_{c}$ from model (1), while Panel C reports year-specific event-study coefficients from model (3). The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below the graph, the DiD coefficient from static versions of models (1) and (3), its standard error, and the mean of the control group in the post-treatment period are reported.
@@
Second, we examine the effect of job loss on donations to political candidates. In Brazil, individuals are allowed to donate up to 10% of their annual income to a political campaign. For candidates, individual contributions serve as an important source for their campaign spending. Consequently, political candidates frequently reward their supporters by offering jobs in the public sector (Colonnelli et al., 2020). Our findings reveal a positive effect on campaign donations in the two cycles after job loss. [^28] Data on campaign contributions is also obtained from the Tribunal Superior Eleitoral. We exclude contributions to individuals’ own campaigns to ensure that the results are not only driven by own candidacies. The estimates are both highly significant and substantial in magnitude. The DiD coefficient indicates a 0.049 percentage point increase in individuals’ probability of donating to a political campaign, corresponding to an 18% increase relative to the control mean. The results suggest that despite experiencing financial setbacks from job layoffs, individuals are not deterred from participating in political activities that require financial contributions. This also aligns with the observed positive impact on party memberships and candidacies, both of which may entail some financial investments in the form of membership fees or campaign funding.
@@
Third, we study the possibility of working for a political campaign. Bazzi and Labanca (2023) show that working for a winning mayoral campaign in Brazil yields substantial private income gains from improved access to jobs in the local bureaucracy. Following their approach, we identify dedicated salaried campaign staff in the RAIS data using information on the industry and legal nature of the employers. [^29] Specifically, we define campaign workers as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Our results show an increase in individuals’ probability of working for a political campaign after job loss. Despite also being a rare event in our sample, the estimates are highly significant and suggest roughly a doubling in the likelihood of becoming a campaign worker. The effects are also highly persistent up to seven years after the layoff which is in line with our findings on party memberships and councilor candidacies.
@@ 4.3 Individual and local heterogeneity
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) for different subgroups of individuals. For education, we differentiate no degree, elementary school, or middle school ($<$ 12 years), high school (12 years), and university education ($>$ 12 years). For age, the sample is split by terciles. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@
Beyond the overall positive effects, Figure 5 reveals substantial differences across education levels, with larger effects on party membership and candidacy among more educated individuals. In quantitative terms, we estimate that individuals with a university degree ($>$ 12 years of education) experience an increase of about 25% in the probabilities of both joining a political party and running for office, while for individuals without a high school degree ($<$ 12 years of education), the estimated effects on both outcomes are below 10%. These patterns are consistent with the opportunity-cost framework presented earlier. Highly educated individuals tend to have better earnings prospects in the private sector, so being displaced in a mass layoff may represent a larger shock to their labor income. Furthermore, the benefits from political engagement – whether in terms of prestige, perceived chances of winning, or future job opportunities – may be higher for more educated individuals, making a political career more attractive. To shed light on these mechanisms, Appendix Figure B.5 examines post-layoff earnings and employment trajectories by education level. The results show that highly-educated individuals indeed suffer substantially larger total earnings losses after the layoff. While their overall employment rate does not improve more rapidly, they are more likely to transition into public-sector jobs, suggesting that political connections may offer them greater potential returns through improved access to such positions.
@@
In terms of the other two individual characteristics considered – age and gender – we do not observe a clear gradient in the effects on new party memberships. In contrast, the impact on candidacies is more pronounced among younger individuals and women. These patterns may reflect differences in labor market attachment and career flexibility. Younger workers often face less entrenched career paths, which can lower the barriers to start a political position after a layoff. [^30] The larger effects for young workers are also consistent with the “impressionable years” hypothesis according to which political attitudes are more malleable at younger ages (e.g., Krosnick and Alwin, 1989). Similarly, given persistent gender gaps in labor market opportunities, women may also face fewer opportunity costs when shifting toward a political career, contributing to their higher likelihood of running for office following displacement.
@@ 4.4 Monetary incentives
Predicted income loss. We start by examining whether predicted income losses due to the layoff are related to the effects of job loss on political entry. [^31] We focus on predicted, rather than actual earnings loss, since the latter may be directly affected by individuals’ behavior after the layoff, such as their decision to run for office. Hilger (2016) and Britto et al. (2022a) conduct similar exercises for the effect of parental job loss on children’s outcomes. The income effects of job loss are predicted as follows. Similar to Schmieder et al. (2023), we first construct for each laid-off worker in our sample an individual-specific measure of the relative earnings losses after layoff by exploiting the fact that we have matched each treated worker to a similar control worker:
@@
Councilor wages. Since 2004, a constitutional amendment has capped councilor salaries based on the population size of the municipality. In municipalities with less than 10,000 inhabitants, councilors can receive a maximum of 20% of state deputy salaries. This share increases for larger municipalities, up to 75% in municipalities with more than 500,000 inhabitants. State deputy salaries, in turn, can get at most 75% of the federal deputy salary. Appendix Table B.4 reports the maximum allowed councilor salary for each population size group, as estimated by Ferraz and Finan (2011). In absolute terms, maximum councilor wages vary between 1,927 BRL (391 USD) and 7,227 BRL (1,466 USD) per month, demonstrating that there is substantial local variation in the financial incentives to run for local councilor.
@@ 4.5 Party disaggregations
(1) (2) (3) (4) (5) (6) (7) (8) All Ideology Local incumbency Union affiliation Left Center Right Yes No Yes No [A] Outcome: new party membership (%) $\delta_{DiD}$ 0.175∗∗∗ 0.045∗∗∗ 0.039∗∗∗ 0.091∗∗∗ 0.024∗∗∗ 0.151∗∗∗ 0.033∗∗∗ 0.051∗∗∗ (0.018) (0.011) (0.010) (0.011) (0.006) (0.018) (0.011) (0.010) Control mean 1.407 0.537 0.353 0.517 0.099 1.308 0.442 0.448 Relative effect 12.4% 8.4% 11.0% 17.6% 24.0% 11.6% 7.5% 11.3% Observations 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 [B] Outcome: candidacy (%) $\delta_{DiD}$ 0.041∗∗∗ 0.012∗∗∗ 0.010∗∗∗ 0.019∗∗∗ 0.005∗∗∗ 0.036∗∗∗ 0.010∗∗∗ 0.012∗∗∗ (0.006) (0.003) (0.003) (0.004) (0.001) (0.006) (0.003) (0.003) Control mean 0.229 0.078 0.060 0.090 0.013 0.216 0.066 0.072 Relative effect 18.1% 15.3% 17.2% 21.1% 43.1% 16.6% 15.5% 16.7% Observations 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 Note: The table reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B). Column (1) presents the baseline results for all parties. Columns (2) to (4) differentiate left, center, and right parties, following the ideological classification of Colonnelli et al. (2022) which is shown in Appendix Table B.5. Columns (5) and (6) distinguish memberships and candidacies in the party of the elected mayor in the individual’s municipality in 2008 and all other non-incumbent parties. Columns (7) and (8) show results separately for parties historically affiliated with labor unions (PT, PDT, PSB, PCB, PSD, and MDB) and all other left or center parties. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Ideology. Layoffs may lead individuals to align with parties whose political programs better address workers’ revised economic interests, for example, through redistributive policies. Moreover, even in the absence of economic platforms, extreme parties may be more successful in exploiting individuals’ resentment of the establishment following economic downturns (e.g., Algan et al., 2017; Autor et al., 2020). To analyze whether layoffs lead to ideological shifts, we distinguish left, center, and right parties, following the classification of Colonnelli et al. (2022). [^32] The ideological classification of parties is shown in Appendix Table B.5. Our results, shown in columns (2) to (4) of Table 2, indicate that job loss leads to an increase in party memberships and candidacies across the ideological spectrum. Relative to the control mean, we find an increase of 8.4%, 11.0% and 17.6% in the likelihood of individuals newly affiliating with left, center and right parties, respectively. The effects on candidacy reveal a similar pattern, with likelihood increases of 15.1%, 17.3%, and 21.3% of running for office in a left, center, or right-wing party. Although we observe larger effects for right-wing parties, our results show that center and left-wing parties also receive more members and candidates following layoffs. Interestingly, the event-study results presented in Appendix Figure B.6 show that the higher increase in affiliations with right-wing parties is driven only by the 2020 cycle, the first local election after Bolsonaro took power in 2019. The rise in right-wing support following economic hardships, which is well documented in the literature (e.g., Colantone and Stanig, 2018; Dippel et al., 2022; Barros and Santos Silva, 2025), appears to be contingent on the current political climate.
@@
Local incumbency. Next, we examine whether individuals punish the party of the local incumbent government after being laid off. Retrospective voting has been widely documented to rationalize the behavior of voters following economic shocks. In particular, individuals may blame incumbent politicians for local economic policies that result in mass layoffs, or for insufficient support policies that mitigate economic hardship after a layoff (e.g., Tilley et al., 2018; Ahlquist et al., 2020). In columns (5) and (6) of Table 2, we compare the effects for the party of the elected mayor in each individual’s municipality before the layoff in 2008 and all other non-incumbent parties. The results show that job loss leads to increased participation in both incumbent and non-incumbent parties. Relative to the control mean, the effects are considerably stronger for incumbent parties, with an increase of 24.0% in the probability of being affiliated with the party of the incumbent and an increase of 43.7% in the likelihood of running for office in the party of the incumbent (as opposed to 11.6% and 16.7% in non-incumbent parties). The results do not support incumbent punishment as a main driver of our results. Instead, running under the incumbent party may raise the likelihood of electoral success, thereby increasing the expected benefits of participation (consistent with our theoretical framework in Section 2.3).
@@
Union affiliation. Lastly, we investigate whether the effects depend on parties’ alignment with labor unions. Unions can influence employees’ political behavior through various activities, such as educating, informing, and mobilizing their members, endorsing selected candidates and policies, and organizing get-out-the-vote campaigns (e.g., Kim and Margalit, 2017; Matzat and Schmeißer, 2023). In Brazil, labor unions have historically been linked to left and centrist parties and have played an important role in advocating for worker interests in domestic politics. Until 2017, unions were required by law to participate in negotiations to formalize dismissals in the context of mass layoffs. [^33] In 2017, the Brazilian labor reform (Law 13,467/17) established that it was no longer mandatory for unions to take part in collective bargaining agreements during mass layoffs. We define union-affiliated parties following the classification of Ogeda et al. (2025) and compare them to all other left-wing or center parties that have no historical ties with labor unions in Brazil. [^34] Union-affiliated parties are PT, PDT, PSB, PCB, PSD, and MDB, while non-union left and center parties are CIDADANIA, PCDOB, PV, PMN, PSOL, PSTU, PCO, SD, PROS, PPL, PMB, REDE, UP, PSDB, PTB, and AVANTE. Our results show an increase in the probability of individuals becoming members or local candidates in both union-affiliated and non-union-affiliated parties (columns (7) and (8) of Table 2). In quantitative terms, we estimate a 7.5% (11.3%) increase in the probability of newly affiliating with union-aligned (non-union-aligned) parties and a 15.4% (16.7%) increase in candidacy probabilities in union-affiliated (non-union-affiliated) parties following mass layoffs. Thus, we find no evidence that unions play a major role in politically mobilizing laid-off individuals.

### v4

@@ 4.2 Main results
We begin by discussing the effects of job loss on labor market trajectories. Figure 1 displays yearly event-study estimates based on model (3). In the year of the layoff, workers experience a sharp decline in labor earnings of around 10,000 BRL (2,028 USD), which correspondents to a 45% reduction relative to baseline earnings in the year prior to the layoff. Moreover, the probability to be formally employed at the end of the year falls by about 60 percentage points. Starting two years after the layoff, both earnings and employment start to continuously recover; yet, the recovery is only partial and does not return to pre-layoff levels. Seven years after the layoff – the longest horizon we can observe in the RAIS data – dismissed workers still earn about 5,000 BRL (1,014 USD) less per year and are 15 percentage points less likely to be formally employed compared to the control group. Appendix Figure B.1 further shows that job loss reduces subsequent job stability, as reflected in a higher frequency of job separations. These findings align with the job loss literature that has widely documented the long-term scarring effects of job displacements (e.g., jacobson1993; eliason2006; davis2011; lachowska2020; schmieder2023). [^29] Britto2022 also show, using survey data on Brazilian workers, that accounting for income from the informal sector reduces the estimated earnings losses following layoffs by only about 10%.
@@
Effect timing. Next, we take a closer look at the dynamics in the estimated treatment effects. Given that deciding to become a candidate and preparing for an election campaign requires time and planning, we start by examining whether our results differ by proximity to the next local election. Appendix Figure B.3 shows results when splitting the sample by the layoff year (2009, 2010, or 2011). We find that the positive effects of job loss on party memberships and candidacies are not driven by one specific treatment year, suggesting that temporal proximity to the local election does not matter. This dynamic also supports the external validity of our results. For instance, there may be the concern that political responses following layoffs in 2009 could differ from other layoffs, given the proximity to the 2008 financial crisis. However, our results show similar effects for layoffs occurring in 2010 and 2011, once the crisis had already faded out in Brazil (barbosa2010).
@@
A. Running for mayor B. Campaign donation Refer to caption Refer to caption C. Campaign workRefer to caption Note: The figure shows results for the effect of job loss on the likelihood of running for municipal mayor (Panel A), donating to a political campaign (Panel B), and working for a political campaign (Panel C). Information on mayoral candidates and campaign donations are obtained from the Tribunal Superior Eleitoral. The data on campaign donations is only available starting from the 2004 election cycle. Following bazzi2023, campaign workers are identified in RAIS as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Panels A and B report cycle-specific event-study coefficients $\delta_{c}$ from model (1), while Panel C reports year-specific event-study coefficients from model (3). The vertical bars depict 95% confidence intervals based on standard errors clustered at the individual level. Below the graph, the DiD coefficient from static versions of models (1) and (3), its standard error, and the mean of the control group in the post-treatment period are reported.
@@
Second, we examine the effect of job loss on donations to political candidates. In Brazil, individuals are allowed to donate up to 10% of their annual income to a political campaign. For candidates, individual contributions serve as an important source for their campaign spending. Consequently, political candidates frequently reward their supporters by offering jobs in the public sector (colonnelli2020). Our findings reveal a positive effect on campaign donations in the two cycles after job loss. [^31] Data on campaign contributions is also obtained from the Tribunal Superior Eleitoral. We exclude contributions to individuals’ own campaigns to ensure that the results are not only driven by own candidacies. The estimates are both highly significant and substantial in magnitude. The DiD coefficient indicates a 0.049 percentage point increase in individuals’ probability of donating to a political campaign, corresponding to an 18% increase relative to the control mean. The results suggest that despite experiencing financial setbacks from job layoffs, individuals are not deterred from participating in political activities that require financial contributions. This also aligns with the observed positive impact on party memberships and candidacies, both of which may entail some financial investments in the form of membership fees or campaign funding.
@@
Third, we study the possibility of working for a political campaign. bazzi2023 show that working for a winning mayoral campaign in Brazil yields substantial private income gains from improved access to jobs in the local bureaucracy. Following their approach, we identify dedicated salaried campaign staff in the RAIS data using information on the industry and legal nature of the employers. [^32] Specifically, we define campaign workers as individuals who have a full-time contract at a firm in industry Atividades de Organizações Políticas (CNAE 1.0: 91928, CNAE 2.0: 94928) or with legal nature Partido Político (3123) or Candidato a Cargo Político Eletivo (4090). Our results show an increase in individuals’ probability of working for a political campaign after job loss. Despite also being a rare event in our sample, the estimates are highly significant and suggest roughly a doubling in the likelihood of becoming a campaign worker. The effects are also highly persistent up to seven years after the layoff which is in line with our findings on party memberships and councilor candidacies.
@@ 4.3 Individual and local heterogeneity
A. New party membership B. Candidacy Refer to caption Refer to caption Note: The figure reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B) for different subgroups of individuals. For education, we differentiate no degree, elementary school, or middle school (<< 12 years), high school (12 years), and university education (>> 12 years). For age, the sample is split by terciles. The horizontal bars depict 95% confidence intervals based on standard errors clustered at the individual level. All coefficients and standard errors are scaled by the group-specific mean of the control group in the post-treatment period (average across 2012, 2016, and 2020).
@@
Beyond the overall positive effects, Figure 5 reveals substantial differences across education levels, with larger effects on party membership and candidacy among more educated individuals. In quantitative terms, we estimate that individuals with a university degree (>> 12 years of education) experience an increase of about 25% in the probabilities of both joining a political party and running for office, while for individuals without a high school degree (<< 12 years of education), the estimated effects on both outcomes are below 10%. These patterns are consistent with the opportunity-cost framework presented earlier. Highly educated individuals tend to have better earnings prospects in the private sector, so being displaced in a mass layoff may represent a larger shock to their labor income. Furthermore, the benefits from political engagement – whether in terms of prestige, perceived chances of winning, or future job opportunities – may be higher for more educated individuals, making a political career more attractive. To shed light on these mechanisms, Appendix Figure B.5 examines post-layoff earnings and employment trajectories by education level. The results show that highly-educated individuals indeed suffer substantially larger total earnings losses after the layoff. While their overall employment rate does not improve more rapidly, they are more likely to transition into public-sector jobs, suggesting that political connections may offer them greater potential returns through improved access to such positions.
@@
In terms of the other two individual characteristics considered – age and gender – we do not observe a clear gradient in the effects on new party memberships. In contrast, the impact on candidacies is more pronounced among younger individuals and women. These patterns may reflect differences in labor market attachment and career flexibility. Younger workers often face less entrenched career paths, which can lower the barriers to start a political position after a layoff. [^33] The larger effects for young workers are also consistent with the “impressionable years” hypothesis according to which political attitudes are more malleable at younger ages (e.g., krosnick1989). Similarly, given persistent gender gaps in labor market opportunities, women may also face fewer opportunity costs when shifting toward a political career, contributing to their higher likelihood of running for office following displacement.
@@ 4.4 Monetary incentives
Predicted income loss. We start by examining whether predicted income losses due to the layoff are related to the effects of job loss on political entry. [^34] We focus on predicted, rather than actual earnings loss, since the latter may be directly affected by individuals’ behavior after the layoff, such as their decision to run for office. hilger2016 and Britto2022a conduct similar exercises for the effect of parental job loss on children’s outcomes. The income effects of job loss are predicted as follows. Similar to schmieder2023, we first construct for each laid-off worker in our sample an individual-specific measure of the relative earnings losses after layoff by exploiting the fact that we have matched each treated worker to a similar control worker:
@@
Councilor wages. Since 2004, a constitutional amendment has capped councilor salaries based on the population size of the municipality. In municipalities with less than 10,000 inhabitants, councilors can receive a maximum of 20% of state deputy salaries. This share increases for larger municipalities, up to 75% in municipalities with more than 500,000 inhabitants. State deputy salaries, in turn, can get at most 75% of the federal deputy salary. Appendix Table B.4 reports the maximum allowed councilor salary for each population size group, as estimated by Ferraz2011. In absolute terms, maximum councilor wages vary between 1,927 BRL (391 USD) and 7,227 BRL (1,466 USD) per month, demonstrating that there is substantial local variation in the financial incentives to run for local councilor.
@@ 4.5 Party disaggregations
(1) (2) (3) (4) (5) (6) (7) (8) All Ideology Local incumbency Union affiliation Left Center Right Yes No Yes No [A] Outcome: new party membership (%) $\delta_{DiD}$ 0.175∗∗∗ 0.045∗∗∗ 0.039∗∗∗ 0.091∗∗∗ 0.024∗∗∗ 0.151∗∗∗ 0.033∗∗∗ 0.051∗∗∗ (0.018) (0.011) (0.010) (0.011) (0.006) (0.018) (0.011) (0.010) Control mean 1.407 0.537 0.353 0.517 0.099 1.308 0.442 0.448 Relative effect 12.4% 8.4% 11.0% 17.6% 24.0% 11.6% 7.5% 11.3% Observations 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 [B] Outcome: candidacy (%) $\delta_{DiD}$ 0.041∗∗∗ 0.012∗∗∗ 0.010∗∗∗ 0.019∗∗∗ 0.005∗∗∗ 0.036∗∗∗ 0.010∗∗∗ 0.012∗∗∗ (0.006) (0.003) (0.003) (0.004) (0.001) (0.006) (0.003) (0.003) Control mean 0.229 0.078 0.060 0.090 0.013 0.216 0.066 0.072 Relative effect 18.1% 15.3% 17.2% 21.1% 43.1% 16.6% 15.5% 16.7% Observations 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 11,317,632 Note: The table reports DiD coefficients, estimated in model (2), for the effect of job loss on the likelihood of newly registering as a party member (Panel A) and running for local councilor (Panel B). Column (1) presents the baseline results for all parties. Columns (2) to (4) differentiate left, center, and right parties, following the ideological classification of Colonnelli2022 which is shown in Appendix Table B.5. Columns (5) and (6) distinguish memberships and candidacies in the party of the elected mayor in the individual’s municipality in 2008 and all other non-incumbent parties. Columns (7) and (8) show results separately for parties historically affiliated with labor unions (PT, PDT, PSB, PCB, PSD, and MDB) and all other left or center parties. All coefficients, standard errors, and control means have been scaled by 100, such that effects are interpreted in terms of percentage points. $p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
@@
Ideology. Layoffs may lead individuals to align with parties whose political programs better address workers’ revised economic interests, for example, through redistributive policies. Moreover, even in the absence of economic platforms, extreme parties may be more successful in exploiting individuals’ resentment of the establishment following economic downturns (e.g., Algan2017a; Autor). To analyze whether layoffs lead to ideological shifts, we distinguish left, center, and right parties, following the classification of Colonnelli2022.353535The ideological classification of parties is shown in Appendix Table B.5. Our results, shown in columns (2) to (4) of Table 2, indicate that job loss leads to an increase in party memberships and candidacies across the ideological spectrum. Relative to the control mean, we find an increase of 8.4%, 11.0% and 17.6% in the likelihood of individuals newly affiliating with left, center and right parties, respectively. The effects on candidacy reveal a similar pattern, with likelihood increases of 15.1%, 17.3%, and 21.3% of running for office in a left, center, or right-wing party. Although we observe larger effects for right-wing parties, our results show that center and left-wing parties also receive more members and candidates following layoffs. Interestingly, the event-study results presented in Appendix Figure B.6 show that the higher increase in affiliations with right-wing parties is driven only by the 2020 cycle, the first local election after Bolsonaro took power in 2019. The rise in right-wing support following economic hardships, which is well documented in the literature (e.g., colantone2018; dippel2022; Barros2019), appears to be contingent on the current political climate.
@@
Local incumbency. Next, we examine whether individuals punish the party of the local incumbent government after being laid off. Retrospective voting has been widely documented to rationalize the behavior of voters following economic shocks. In particular, individuals may blame incumbent politicians for local economic policies that result in mass layoffs, or for insufficient support policies that mitigate economic hardship after a layoff (e.g., Tilley2018; Ahlquist2020). In columns (5) and (6) of Table 2, we compare the effects for the party of the elected mayor in each individual’s municipality before the layoff in 2008 and all other non-incumbent parties. The results show that job loss leads to increased participation in both incumbent and non-incumbent parties. Relative to the control mean, the effects are considerably stronger for incumbent parties, with an increase of 24.0% in the probability of being affiliated with the party of the incumbent and an increase of 43.7% in the likelihood of running for office in the party of the incumbent (as opposed to 11.6% and 16.7% in non-incumbent parties). The results do not support incumbent punishment as a main driver of our results. Instead, running under the incumbent party may raise the likelihood of electoral success, thereby increasing the expected benefits of participation (consistent with our theoretical framework in Section 2.3).
@@
Union affiliation. Lastly, we investigate whether the effects depend on parties’ alignment with labor unions. Unions can influence employees’ political behavior through various activities, such as educating, informing, and mobilizing their members, endorsing selected candidates and policies, and organizing get-out-the-vote campaigns (e.g., kim2017; matzat2023). In Brazil, labor unions have historically been linked to left and centrist parties and have played an important role in advocating for worker interests in domestic politics. Until 2017, unions were required by law to participate in negotiations to formalize dismissals in the context of mass layoffs. [^36] In 2017, the Brazilian labor reform (Law 13,467/17) established that it was no longer mandatory for unions to take part in collective bargaining agreements during mass layoffs. We define union-affiliated parties following the classification of ogeda2021 and compare them to all other left-wing or center parties that have no historical ties with labor unions in Brazil. [^37] Union-affiliated parties are PT, PDT, PSB, PCB, PSD, and MDB, while non-union left and center parties are CIDADANIA, PCDOB, PV, PMN, PSOL, PSTU, PCO, SD, PROS, PPL, PMB, REDE, UP, PSDB, PTB, and AVANTE. Our results show an increase in the probability of individuals becoming members or local candidates in both union-affiliated and non-union-affiliated parties (columns (7) and (8) of Table 2). In quantitative terms, we estimate a 7.5% (11.3%) increase in the probability of newly affiliating with union-aligned (non-union-aligned) parties and a 15.4% (16.7%) increase in candidacy probabilities in union-affiliated (non-union-affiliated) parties following mass layoffs. Thus, we find no evidence that unions play a major role in politically mobilizing laid-off individuals.